Please help with 1, 3, and 4 :)

A. ΔHsolution of NaOH(s) Phase Change:
Mass of NaOH: 1.81 g
ΔT for reaction A : 6 °C

B. ΔHsolution for neutralization of HCl(aq) and NaOH(aq):
Volume of HCl: .05 L
Volume of NaOH: .05 L
Volume total: .10 L
Molarity of HCl: 2.08 M
Molarity of NaOH: 2.08 M
ΔT for reaction B: 12°C

C. ΔH for Reaction of HCl(aq) and NaOH(s):
Volume of HCl: .055 L
Volume of water: .045 L
Volume total: .1 L
Molarity of HCl: 2.08 M
Mass of NaOH: 4.3 g
ΔT for reaction C: 22°C

D. ΔH for Solution of KCl(s) in water:
ΔT for reaction D: -3.5°C
Mass of KCl: 5 g
Volume of water: .05 LB. AH for neutralization of HClaq and NaOHaq: Calculate the molarity of the resulting sodium chloride solution. 2.08 Submit A
Answer #1

SOLN B

NaOH: moles = molarity * volume

2.08 M x (0.05 L) = 0.104 moles NaOH

same is the mole of HCl

hence both will neutralize each other

hence moles of NaCl forme = 0.104 moles

Molarity = moles / total volume = 0.104 / 0.1 = 1.04 M

Heat = mass x T change x specific heat

mass = moles * molar mass = 0.104 *40 = 4.16 g

Heat = 4.16 g x 12 C x 1.00 cal/gC = 49.92 cal

dH = heat / mol = 49.92 cal / 0.104 mol = 480 cal/mol (of HCl or NaOH)

1cal = 4.184 J

hence 480 cal = 2007 J or 2.007 kJ/mol